# DPO Meets PPO: Reinforced Token Optimization for RLHF

### v1

@@ 3.3 Advantages of Token-Wise MDP over Sentence-Wise Bandit
Figure 2: An illustration of our efficient learning algorithm for the token-wise reward setting with $A=2$, $H=3$, and $\xi=1$. Here * and $\dagger$ represent real numbers between 0 and 1/8. We do not specify their exact values as they do not influence the optimal path. All nodes in $\mathcal{N}$ are colored red, while other nodes are blue, with the optimal leaf node $1/2$ emphasized in dark blue. Each node $y_{1:h}$ is labelled with $\pi^{*}(y_{1:h}{\,|\,}x)$. If a non-optimal path (response) is selected, one red node in $\mathcal{N}$ will be identified and added to $\mathcal{N}^{\prime}$, and all paths containing this node will be deleted. Here we visualize the process of choosing a path ending with *, $\dagger$, and $1/4$, respectively. At most $A^{\min\{\zeta+1,H\}}=4$ samples are needed to identify the optimal response.
@@
If only the sentence-level reward $r_{s}$ is available, the learner must try every possible response and determine the optimal one by ranking the collected sentence-level reward signals, resulting in a sample complexity of $A^{H}$. Instead, we consider a binary tree with depth $H+1$, where each node is indexed by some token sequence $y_{1:h}$ and has $A$ children $\{(y_{1:h},y_{h+1})\}_{y_{h+1}\in\mathcal{A}}$. All $A^{H}$ leaf nodes denote a unique prompt-response pair $(x,y_{1:H})$. We define a set of nodes as
$\mathcal{N}=\big{\{}y_{1:h}:\pi^{*}(y_{1:h}{\,|\,}x)<A^{-\xi},\pi^{*}(y_{1:h-1}{\,|\,}x)\geq A^{-\xi}\big{\}}.$ (3.7)
Our key observation is that $|\mathcal{N}|\leq A^{\xi+1}$. We also maintain a node set $\mathcal{N}^{\prime}$. Initially, we set $\mathcal{N}^{\prime}=\emptyset$. If $\mathcal{N}^{\prime}$ is updated, we delete all paths containing some node in $\mathcal{N}^{\prime}$. Each query of a new path (response with length $H$) will identify an additional node in $\mathcal{N}$. Then we add it to $\mathcal{N}^{\prime}$ and delete all paths containing some node in $\mathcal{N}^{\prime}$. This operation ends after at most $A^{\xi+1}$ iterations. Finally, ranking all gathered rewards identifies the optimal $y^{*}=y_{1:H}^{*}$. Together with the fact that there exists as most $A^{H}$ nodes, we finish the proof of Proposition 3.2. To facilitate understanding, we visualize a simplified learning process in Figure 2. ∎
@@ 5.1 Experimental Setups
We study the performance of our model on the single-turn dialogue generation task (Bai et al., 2022). Given a text sequence ($x$) representing dialogue history between the user and the assistant, the goal of the task is to generate a helpful response ($y$) as the answer. For this purpose, we utilize the helpful subset of the Anthropic Helpful and Harmless (HH-RLHF) dialogue dataset [^3] https://huggingface.co/datasets/Anthropic/hh-rlhf (Bai et al., 2022). Each sample of the HH-RLHF dataset is accompanied by a history and two alternative responses, with preferences annotated by humans. We provide an example of the HH-RLHF dataset in Appendix D.1.
@@
We use two metrics to evaluate the alignment performance of different methods: oracle reward evaluation and GPT-4 evaluation. For oracle reward evaluation, we employ an open-sourced reward model  [^4] https://huggingface.co/weqweasdas/RM-Mistral-7B as oracle, which is trained from Mistral-7B (Jiang et al., 2023) and achieves one of the highest accuracy on the Anthropic Helpful and Harmless dialogue task (Bai et al., 2022). For each pair of models, given the same prompt, we generate the responses using both models and calculate the rewards of the responses by the oracle reward model. We then compare the rewards of the two models and report the win rates between them. The GPT-4 evaluation, on the other hand, leverages the capabilities of GPT-4 itself and has been demonstrated to correlate with human evaluations (Rafailov et al., 2023) well. Given the two responses for the same prompt using two models, we ask GPT-4 about which one is better and calculate the win rates, following Rafailov et al. (2023). The prompt for GPT-4 evaluation is provided in Table 6 in Appendix D.3. For each evaluation by the oracle reward model, 400 dialogue histories from the test dataset are sampled, and for each evaluation by GPT-4, 100 dialogue histories are sampled.
@@ 5.2 Experimental Results
The experiment results of our proposed method and baselines are detailed in Table 1. This table meticulously presents the win rates between different models, assessed through both the oracle reward and the GPT-4.
@@
RTO 0.500 0.556 0.629 0.600 0.611
DPO 0.444 0.500 0.578 0.596 0.573
SFT 0.371 0.422 0.500 0.487 0.525
PPO 0.400 0.404 0.513 0.500 0.511
DPPO 0.389 0.427 0.475 0.489 0.500
Table 1(a): Win rate evaluated by oracle reward model.
@@
RTO 0.50 0.51 0.56 0.61 0.56
DPO 0.49 0.50 0.52 0.63 0.51
SFT 0.44 0.48 0.50 0.58 0.46
PPO 0.39 0.37 0.42 0.50 0.45
DPPO 0.44 0.49 0.54 0.55 0.50
Table 1(b): Win rate evaluated by GPT-4
Table 1: The left table shows the win rates between each pair of models evaluated by the oracle reward evaluation, and the right one shows the win rates evaluated by GPT-4. The value in line $i$ column $j$ represents the win rate of the model in row $i$ against the model in column $j$. We can see that the model trained by RTO achieves win rates over $50\%$ against all other baselines.
From these results, we can see that the model trained by RTO achieves win rates over $50\%$ against all other baselines, especially compared to DPO, evaluated by both the oracle reward model and GPT-4. This highlights the effectiveness of the RTO algorithm in alignment tasks. Furthermore, the model trained by RTO gets a win rate of $61.1\%$ evaluated by the oracle reward model and a win rate of $56\%$ evaluated by GPT-4 over the DPPO algorithm. This implies that the token-wise reward mechanism significantly improves the performance of the RL algorithm in training models. To further investigate the benefits of the token-wise reward mechanism in the optimization process, we compare the estimated reward during the training period in Figure 3. In this figure, the x-axis represents the training iterations (1 epoch roughly corresponds to 160 PPO training iterations). The y-axis represents the reward given by the implicit reward model derived from the DPO model (the reward model used in training) per batch. As we can see, in one epoch, the reward of the model trained by RTO can achieve about $0.1$, while the reward of the model trained by DPPO is roughly $-0.25$. The results demonstrate that the token-wise reward mechanism significantly enhances the training process, leading to a remarkably higher reward. All these empirical findings demonstrate the token-wise reward mechanism’s advantage in improving model performance.
Figure 3: The reward curve of DPPO and RTO during training. The reward is given by the implicit reward model $\beta\log\frac{\pi_{\mathrm{dpo}}(y{\,|\,}x)}{\pi_{\mathrm{ref}}(x{\,|\,}y)}$ optimized by DPO. The x-axis represents the training steps, and the y-axis represents the reward values.
@@ C.2 Autoregressive Policy
Appendix D Additional Experimental Details
D.1 A Sample of the HH-RLHF Dataset
@@ D.2 Training Configurations
We provide the training configuration of SFT, DPO, PPO, DPPO, and RTO below. In the table of the training configuration of the standard PPO algorithm, we also present the configuration of training the reward model used in the PPO algorithm in this table.
@@
Init KL Coefficient ($\beta$) 0.05
@@
Learning Rate (Reward Model) 1e-5
@@ D.3 Evaluation Details
We employ the implicit reward model, derived from the DPO model, as the Oracle reward model. This evaluation is conducted for both dialogue generation and summarization tasks. To assess the performance across models, we use top-p sampling and set $p=0.99$ and temperature $\tau=0.9$ to generate completions for 400 prompts from the test set. These samples are then compared based on their rewards to calculate the win rate of one model over another.

### v2

@@ 3.3 Advantages of Token-Wise MDP over Sentence-Wise Bandit
Figure 2: An illustration of our efficient learning algorithm for the token-wise reward setting with $A=2$, $H=3$, and $\xi=1$. Here * and $\dagger$ represent real numbers between 0 and 1/8. We do not specify their exact values as they do not influence the optimal path. All nodes in $\mathcal{N}$ are colored red, while other nodes are blue, with the optimal leaf node $1/2$ emphasized in dark blue. Each node $y_{1:h}$ is labelled with $\pi^{*}(y_{1:h}{\,|\,}x)$. If a non-optimal path (response) is selected, one red node in $\mathcal{N}$ will be identified, and all paths containing this node will be deleted. Here we visualize the process of choosing a path ending with *, $\dagger$, and $1/4$, respectively. At most $A^{\min\{\xi+1,H\}}=4$ samples are needed to identify the optimal response.
@@
If only the sentence-level reward $r_{s}$ is available, the learner must try every possible response and determine the optimal one by ranking the collected sentence-level reward signals, resulting in a sample complexity of $A^{H}$. Instead, we consider a binary tree with depth $H+1$, where each node is indexed by some token sequence $y_{1:h}$ and has $A$ children $\{(y_{1:h},y_{h+1})\}_{y_{h+1}\in\mathcal{A}}$. All $A^{H}$ leaf nodes denote a unique prompt-response pair $(x,y_{1:H})$. We define two disjoint node sets:
$\mathcal{N}=\big{\{}y_{1:h}:\pi^{*}(y_{1:h}{\,|\,}x)<A^{-\xi},\pi^{*}(y_{1:h-1}{\,|\,}x)\geq A^{-\xi}\big{\}},\qquad\mathcal{N}^{*}\big{\{}y_{1:H}:\pi^{*}(y_{1:H}{\,|\,}x)\geq A^{-\xi}\big{\}}.$ (3.7)
Our key observations are that (i) each path must contain a node in $\mathcal{N}$ or $\mathcal{N}^{*}$, (ii) the path containing the node in $\mathcal{N}$ is suboptimal; and (iii) $|\mathcal{N}\cup\mathcal{N}^{*}|\leq A^{\xi+1}$. The exploration strategy is to query a new path that does not contain the nodes in $\mathcal{N}\cup\mathcal{N}^{\prime}$ that have been visited. Since each query of a new path (response with length $H$) can identify a new additional node in $\mathcal{N}\cup\mathcal{N}^{*}$, after at least $A^{\xi+1}$ queries, we collect a set of paths where each node in $\mathcal{N}\cup\mathcal{N}^{*}$ belongs to one of the paths. Finally, ranking all gathered rewards of the node in $\mathcal{N}^{*}$ identifies the optimal $y^{*}=y_{1:H}^{*}$. Together with the fact that there exists as most $A^{H}$ nodes, we finish the proof of Theorem 3.2. To facilitate understanding, we visualize a simplified learning process in Figure 2. ∎
@@ 5.1 Experimental Setups
We consider two text generation tasks, including the Dialogue Generation (Bai et al., 2022) and Summarization task (Völske et al., 2017). Given a text sequence ($x$) representing dialogue history between the user and the assistant, the goal of the Dialogue Generation task is to generate a helpful response ($y$) as the answer. For this purpose, we utilize the helpful subset of the Anthropic Helpful and Harmless (HH-RLHF) dialogue dataset [^3] https://huggingface.co/datasets/Anthropic/hh-rlhf (Bai et al., 2022). Each sample of the HH-RLHF dataset is accompanied by a history and two alternative responses, with preferences annotated by humans. For the Summarization task, the model is required to generate a concise summary for a given post from the Reddit forum. Specifically, we self-supervise fine-tune the foundational model using the Reddit TL;DR summarization dataset (Völske et al., 2017), where each data point comprises a post $x$ and its corresponding summary $y$. Subsequently, we align the model with human preferences using its preference version [^4] https://huggingface.co/datasets/openai/summarize_from_feedback, where each data point comprises a post and two summaries, with preferences annotated by humans. To facilitate readers, we provide examples of the HH-RLHF and TL;DR datasets in Appendix D.1.
@@
We primarily assess the alignment performance of various methods using GPT-4. The GPT-4 evaluation harnesses the capabilities of GPT-4 itself and has been shown to align well with human evaluations (Rafailov et al., 2023). For the same prompt, we provide GPT-4 with two responses generated by two different models and ask it to determine which one is superior. We then calculate the win rates, following Rafailov et al. (2023). The prompts for GPT-4 evaluation of both HH-RLHF and TL;DR are presented in Appendix D.3. For each GPT-4 evaluation, we use 100 samples. In addition to GPT-4 evaluation, we also employ oracle reward evaluation for HH-RLHF. For this, we employ an open-sourced reward model  [^5] https://huggingface.co/weqweasdas/RM-Mistral-7B as the oracle, which is trained from Mistral-7B (Jiang et al., 2023) and achieves one of the highest accuracy on the Anthropic Helpful and Harmless dialogue task (Bai et al., 2022). Given 400 prompts, we generate the responses using all baselines and calculate the average rewards of the responses by the oracle reward model.
@@ 5.2 Experimental Results
Win Rates.
The experiment results of our proposed method and baselines on HH-RLHF are detailed in Table 1(a) and Figure 3(c). This table meticulously presents the win rates between different models, assessed through both the oracle model and the GPT-4.
@@
RTO 0.50 0.51 0.56 0.52 0.56
DPO 0.49 0.50 0.52 0.54 0.51
SFT 0.44 0.48 0.50 0.41 0.46
PPO 0.48 0.46 0.59 0.50 0.52
DPPO 0.44 0.49 0.54 0.48 0.50
Table 1(a): HH-RLHF win rates evaluated by GPT-4.
@@
RTO 0.50 0.61 0.67 0.67 0.67
DPO 0.39 0.50 0.58 0.59 0.50
SFT 0.33 0.42 0.50 0.59 0.49
PPO 0.33 0.41 0.41 0.50 0.40
DPPO 0.33 0.50 0.51 0.60 0.50
Table 1(b): TL;DR win rates evaluated by GPT-4
Table 1: Win rates between each pair of models evaluated by GPT-4. The value in line $i$ column $j$ represents the win rate of the model in row $i$ against the model in column $j$.
((a)) HH-RLHF Evaluated by GPT-4
((b)) TL;DR Evaluated by GPT-4
((c)) HH-RLHF Evaluated by oracle model
Figure 3: Subfigures (a) and (b) show the win rates of RTO across different sampling temperatures, while Subfigure (c) visualizes the mean reward of different baselines evaluated by the oracle model.
From these results, we can see that the model trained by RTO achieves win rates over $50\%$ against all other baselines, especially compared to DPO, evaluated by both the oracle reward model and GPT-4. This highlights the effectiveness of the RTO algorithm in alignment tasks. Furthermore, the model trained by RTO gets a win rate of $61.1\%$ evaluated by the oracle reward model and a win rate of $56\%$ evaluated by GPT-4 over the DPPO algorithm. This implies that the token-wise reward mechanism significantly improves the performance of the RL algorithm in training models. Table 1(b) presents the performance of our method on the TL;DR dataset. We can see that the model trained by RTO outperforms all other baselines. Specifically, we achieve a win rate of $61\%$ over the DPO algorithm evaluated by GPT-4. This illustrates the effectiveness of the RTO algorithm in a real-world text summarization task. All these empirical findings demonstrate the token-wise reward mechanism’s advantage in improving model performance.
Ablation Studies on Temperatures.
We further compare the model trained by RTO to other baselines on both datasets across different temperatures. In Figure 3(c), we present the results for these methods as we vary the temperature. We can observe that the RTO model consistently demonstrates superior performance compared to other baselines, highlighting its robustness across different temperatures.
Optimization Process Curves.
To further investigate the benefits of the token-wise reward mechanism in the optimization process, we compare the estimated reward during the training period in Figure 4 and Figure 4. In both figures, the x-axis represents the training iterations. The y-axis represents the reward given by the implicit reward model derived from the DPO model (the reward model used in training) per batch. As we can see, in one epoch (roughly corresponds to 160 PPO training iterations in Figure 4 and 240 in Figure 4), the reward of the model trained by RTO on HH-RLHF can achieve about $0.1$, while the reward of the model trained by DPPO is roughly $-0.25$; similarly, the reward of the model trained by RTO on TL;DR can achieve about $0.4$, while the reward of the model trained by DPPO is roughly $-0.2$. The results demonstrate that the token-wise reward mechanism significantly enhances the training process, leading to a remarkably higher reward.
Figure 4: The reward curve of DPPO and RTO during training. The reward is given by the implicit reward model $\beta\log\frac{\pi_{\mathrm{dpo}}(y{\,|\,}x)}{\pi_{\mathrm{ref}}(y{\,|\,}x)}$ optimized by DPO. The x-axis represents the training steps, and the y-axis represents the reward values. Figure  4 shows the reward curve of HH-RLHF, and Figure  4 shows the reward curve of TL;DR.
@@ C.2 Autoregressive Policy
Appendix D Experimental Details
D.1 Examples of Datasets
D.1.1 A Sample of the HH-RLHF Dataset
@@
D.1.2 A Sample of the TL;DR Dataset
Prompt: SUBREDDIT: r/AskReddit TITLE: Reddit, what event drove you to cry in the bathroom at work? POST: Yesterday, I finally became that girl who goes into the bathroom to cry while at work. I work at a domestic violence shelter, and normally I’m pretty capable of brushing things off. I’m somewhat ashamed to say that it was not secondhand truama that led me to weep in the bathroom stall like a little girl, but my coworkers. It had been a rough day, which are pretty normal around here, but it was a tolerable level of rougness. My patience was wearing thin and I just wanted to go to the support group for advocates and take a breather. Unfortunately, my coworker decided at that time to demand that I clean one of the recently vacated rooms. Not just clean it, but DEEP clean it. I’m not talking clean-it-like-your-parents-are-coming-home-after-a-weekend-away type clean. I mean, she wanted it hospital-grade clean. She wanted to be able to perform surgery on any surface of that room. The checklist she gave me- handwritten of course- had at least thirty tasks on it. For a dorm-sized room. I lost it, guys. I just completely lost my shit. I told her that I would be happy to help clean that room, but she was absolutely off her rocker if she thought I was going to spend the next four hours cleaning by myself. She was incensed at my apparent refusal, and though I tried to reiterate that I would do it, but not alone, she started screaming for the lead advocate to put me in my place. Well, the lead advocate just didn’t want to deal with the situation and told me to just do it. I was absolutely frustrated, appalled, and overwhelmed. And so…I went into the bathroom and cried. Then I went and cleaned the stupid room. TL;DR: Chosen response:  I was stressed, my lazy coworker demanded I clean every speck of dust from a room alone, I lost my shit, my supervisor sided with my coworker. Rejected response:  Coworker thinks it’s okay to ask me to clean a room she thinks is a dumpster, so I cried. Then I cleaned it.
@@ D.2 Training Configurations
We provide the training configuration of SFT, DPO, PPO, DPPO, and RTO below. In the table of the training configuration of the standard PPO algorithm, we also present the configuration of training the reward model used in the PPO algorithm.
D.2.1 Training Configurations of HH-RLHF
@@
Init KL Coefficient ($\beta$) 0.1
@@
Learning Rate (Reward Model) 5e-5
@@
D.2.2 Training Configurations of TL;DR
SFT
Optimizer AdamW
Learning Rate 1e-5
Batch Size 32
Epochs 1
Table 6: Configurations for supervise fine-tuning.
DPO
Optimizer AdamW
Learning Rate 5e-6
KL Coefficient ($\beta$) 0.1
Batch Size 32
Epochs 1
Table 7: Configurations for DPO.
PPO
Optimizer (PPO) Adam
Optimizer (Reward Model) AdamW
Mini Batch Size in PPO 16
Init KL Coefficient ($\beta$) 0.03
Learning Rate (PPO) 3e-6
Learning Rate (Reward Model) 3e-6
Batch Size Per PPO Iteration 256
Epochs of PPO Update Per Iteration 2
Batch Size (Reward Model) 128
Training Epochs (PPO and Reward Model) 1
Maximum Sequence Length 512
RTO and DPPO
Optimizer Adam
Learning Rate 3e-6
Training Epochs 1
Mini Batch Size in PPO 16
DPO KL Coefficient $\beta_{1}$ 0.1
Init KL Coefficient $\beta_{2}$ (RTO) 0.05
Init KL Coefficient $\beta_{2}$ (DPPO) 0.05
Batch Size Per PPO Iteration 256
Maximum Sequence Length 512
Epochs of PPO Update Per Iteration 2
Table 9: Configurations for RTO and DPPO.
@@ D.3 Evaluation Details
we employ an open-sourced reward model  [^6] https://huggingface.co/weqweasdas/RM-Mistral-7B as oracle, which is trained from Mistral-7B (Jiang et al., 2023) and achieves one of the highest accuracy on the Anthropic Helpful and Harmless dialogue task (Bai et al., 2022). This evaluation is conducted only for dialogue generation tasks. To assess the performance across models, we use top-p sampling and set $p=0.99$ and temperature $\tau=0.9$ to generate completions for 400 prompts from the test set. These samples are then compared based on their rewards to calculate the win rate of one model over another.
@@
Prompt for GPT-4 evaluation in summarization task.
Which of the following summaries does a better job of summarizing the most important points in the given forum post, without including unimportant or irrelevant details? A good summary is both precise and concise.
Post: <the forum post>
Summary A: <either the test method or baseline>
Summary B: <the other summarization>
FIRST provide a one-sentence comparison of the two summaries, explaining which you prefer and why. SECOND, on a new line, state only ”A” or ”B” to indicate your choice. Your response should use the format:
Comparison: <one-sentence comparison and explanation>
Comparison:
Preferred: <"A" or "B">
Table 11: Prompt for GPT-4 evaluation in summarization task.